# DPDFNet: Boosting DeepFilterNet2 via Dual-Path RNN

### v1

@@ I Introduction
Early real-time speech enhancement often combined traditional signal processing with compact neural networks. Valin [rnnoise] proposed RNNoise, which predicts critical-band gains from features such as Bark-frequency cepstral coefficients and pitch correlations, then applies a frequency-domain pitch filter. This hybrid design achieves very low latency and embedded-class efficiency, and it outperforms minimum mean-square spectral estimators in perceptual quality. Research then advanced sequence modeling for streaming noise suppression. Westhausen and Meyer [DTLN] introduced DTLN with two cascaded cores. One core operates in the frequency-domain and the other uses a learned time-domain representation. Both rely on LSTM layers with instant layer normalization, enabling stable frame-synchronous processing with fewer than one million parameters and benefiting from both magnitude and phase cues. Braun and Tashev proposed NSNet2 [nsnet2], a recurrent model that estimates spectral suppression gains from log power spectral features. Its level-invariant loss normalization and broad augmentation over SNR, spectral shape, and signal level yield a robust real-time baseline. Another line of work moved to the waveform domain. Defossez et al. [DEMUCS] proposed DEMUCS, an encoder-decoder with U-Net-style skip connections and a recurrent bottleneck that operates directly on raw audio. Training uses a mix of waveform loss and multi-resolution spectral losses with strong augmentation. This approach improves naturalness and intelligibility under real-time constraints. Kong et al. proposed CleanUNet [CleanUNet] as a refinement of this design. It strengthens the bottleneck with stacked masked self-attention and combines waveform and spectral objectives, including a high-band loss that improves the handling of silence and high-frequency detail. Although relatively large, CleanUNet maintains low latency and delivers consistent quality. Work in the time-frequency domain focused on stronger cross-band modeling while preserving local structure. Hao et al. [Fullsubnet] proposed FullSubNet, which encodes long-range context with a full-band LSTM. The frame-wise embedding is concatenated with local frequency neighborhoods and passed to a shared sub-band LSTM that predicts a complex ideal ratio mask [cIRM] (cIRM) with tanh compression. This design links global spectral trends with local stationarity cues. Dual-path recurrent models were optimized to balance efficiency and streaming performance. Le et al. [DPCRN] proposed DPCRN, an encoder-decoder with causal two-dimensional convolutions and skip connections. A dual-path module models spectral structure within each frame and temporal dynamics across frames, under causal constraints. Rong et al. proposed GTCRN [GTCRN] as a direct follow-up. It uses grouped temporal convolutions and grouped dual-path recurrent units, merges high-frequency bands through an equivalent rectangular bandwidth (ERB) filter bank, and applies temporal recurrent attention. The result keeps the parameter count in the tens of thousands while remaining competitive with heavier systems. Most recently, Pei et al. proposed aTENNuate [aTENNuate], a deep state space autoencoder for streaming enhancement on raw audio. The model stacks causal state space blocks with skip connections and light resampling. It targets online operation with steady latency and compact size, and it trains with objectives in both time and frequency domains. Overall, the field has made a substantial progress from classical-plus-RNN hybrids to highly capable pure end-to-end deep neural-network models, for higher intelligibility and naturalness under challenging real-time constraints.
@@
Although a wide array of methods developed in the recent years, including the aforementioned ones, we based our work over Schröter et al. [DeepFilterNet2]’s DeepFilterNet2 architecture: a compact two-stage architecture that offers a strong balance of efficiency and quality for streaming SE. This architecture provides a favorable trade-off between computational complexity and enhancement quality, enabling real-time, low-latency deployment. Moreover, the architecture permits training at lower sampling rates (e.g., 16 kHz) while deploying at higher rates with negligible performance loss. This leverages the wider availability of low-rate data and reduces the computational load of online augmentations. Motivated by both technical considerations and community interest in advanced speech enhancement, we believe that enhancing the current DeepFilterNet2 architecture will contribute much to the field.
@@
Accordingly, we integrate causal dual-path modules into the DeepFilterNet2 encoder, resulting in DPDFNet. This enhanced design delivers a substantial performance gain while preserving the overall architecture. We also add an over-attenuation loss, used alongside the original multi-resolution loss [DeepFilterNet2], to mitigate rare cases where the model over-suppresses the target speech, and proposed a fine-tuning stage to exposed the models to long-range dependencies. Finally, we curate a new evaluation set focused on low-SNR conditions with realistic noise types (car, pub, office, etc.) with multilingual coverage spanning 12 languages from the Speech-MASSIVE test set [Speech-MASSIVE].
@@ II-C Dual-Path RNN Block
Dual-Path Recurrent Neural Networks, introduced by Luo et al. [DPRNN], address long-context modeling through two alternating recurrent stages that operate along complementary axes of the data: an intra stage that aggregates information within each time frame and an inter stage that propagates information across time.
@@
This dual-path strategy was adopted by DPCRN [DPCRN], which replaces the GRU bottleneck in a U-Net-style encoder-decoder with DPRNN blocks and reports strong performance. In addition, DPCRN appends a position-wise fully connected (FC) layer and Layer Normalization (LN) [layer_normalization] after each DPRNN stage, both intra and inter, in order to refine features and stabilize training. Figure 1a depicts the scheme of the proposed DPRNN block.
@@ III-A Datasets and Augmentations
Many recent SE models, including our baseline DeepFilterNet2, have been trained on the widely adopted Deep Noise Suppression 4 (DNS4) Challenge [dns4] dataset. This full-band (48 kHz) dataset provides 760 hours of clean speech recordings in six languages from 3,230 unique speakers, along with 181 hours of noise comprising 62,000 clips across 150 noise classes. It also includes 248 real and 60,000 synthetic RIRs for generating reverberant signals. Although DNS4 remains a widely adopted dataset for training SE models, the clean speech portion is not entirely free of residual noise, which may constrain the achievable performance of trained models. In contrast, a variety of high-quality, open-source wideband (16 kHz) datasets exist for both clean speech and noise, offering valuable resources for more effective model training. Furthermore, since the DeepFilterNet2 architecture can be adapted to full-band operation via a lightweight postprocessing step with negligible impact on performance, utilizing these larger and higher-quality wideband datasets presents a particularly compelling direction for training. Therefore, we downsampled the DNS4 dataset to wideband and enriched it with 4,000 hours of English read speech from the Multilingual LibriSpeech (MLS) corpus [fbmls2020], a large-scale dataset derived from LibriVox released by Facebook. This addition improves both quality and speaker diversity compared to DNS4 alone. To further ensure the cleanest possible speech signals, all audio was processed through the DPCRN [DPCRN] model. Furthermore, we incorporated noise clips from the MUSAN [musan] and FSD50K [fsd50k] datasets to further diversify the noise conditions. To improve model generalization, a set of online data augmentations is applied during training. Random second-order filtering is used to expose the model to natural spectral colorations, improving robustness across devices and recording environments. Gain perturbations are introduced to prevent overfitting to specific loudness levels and to ensure invariance to overall volume. In addition, reverberation with early reflection targets is applied, enabling the model to suppress late reverberation while preserving the direct path and early reflections that are beneficial for speech intelligibility.
@@ III-B New Evaluation Set
To evaluate and compare different SE models, two test sets are commonly employed: (1) the VoiceBank+DEMAND test set [vctk_demand] and (2) the DNS4 blind test set [dns4]. While these datasets have been widely adopted as benchmarks within the single-channel SE domain, they fail to fully capture the diversity of real-world acoustic conditions. For instance, the VoiceBank+DEMAND test set primarily consists of short clips, typically only a few seconds in duration, recorded at relatively high SNRs. Conversely, the DNS4 blind test set provides longer recordings; however, each clip includes only a single speaker and lacks extended periods of background noise without speech - an element often present in realistic environments. Furthermore, both datasets exhibit limited linguistic diversity, as they consist almost exclusively of English speech.
@@ III-C Loss Functions
We adopt the Multi-Resolution loss from  [DeepFilterNet2]. The enhanced signal $y$ is analyzed with several STFTs, specifically with window sizes $i\in\{5,10,20,40\}$ ms, to form the following loss:
@@
denote the complex spectrograms of the enhanced and clean signals, respectively, with phases $\phi_{Y_{i}}$ and $\phi_{S_{i}}$. The magnitude compression exponent set to $c=0.3$ follow the original DeepFilterNet2 training [DeepFilterNet2].
@@ IV Experimental Results
TABLE I: Intrusive, non-intrusive and PRISM scores on our new multilingual low-SNR test set for open-source causal models and our DPDFNet-$\{k\}$ variants ($k$ = number of DPRNN blocks). The baseline uses the DeepFilterNet2 architecture within our training framework.
@@
Noisy 1.27 83.2 0.38 2.02 1.66 1.56 2.53 2.00 1.69 3.40 2.68 2.68 0.04
DTLN [DTLN] 0.99 0.12 1.94 88.8 10.83 2.51 3.50 2.15 2.85 2.13 2.51 2.79 2.39 2.85 0.46
GTCRN [GTCRN] 0.023 0.039 2.00 87.2 9.11 2.56 3.76 2.24 3.00 2.53 3.27 3.17 2.80 3.01 0.49
RNNoise [rnnoise] 0.087 0.04 2.05 88.5 8.97 2.81 3.93 2.50 3.01 1.88 3.68 2.20 2.19 3.14 0.51
NSNet2 [nsnet2] 2.60 0.26 2.06 86.8 8.63 2.68 3.82 2.36 2.92 2.63 3.39 3.25 2.76 3.30 0.51
FullSubNet [Fullsubnet] 5.60 30.00 2.19 89.6 10.63 2.90 3.41 2.43 3.06 2.76 2.78 3.83 3.41 3.29 0.63
aTENNuate [aTENNuate] 0.80 0.33 2.43 88.9 9.65 3.05 4.03 2.74 2.94 3.01 3.35 3.90 2.97 3.92 0.69
DPCRN [DPCRN] 0.53 1.10 2.47 90.7 11.93 2.73 3.78 2.39 3.03 2.92 3.11 3.61 3.23 3.40 0.69
CleanUNet [CleanUNet] 46.07 15.44 2.38 91.2 11.67 2.97 3.95 2.66 3.07 2.57 3.41 2.72 2.86 3.57 0.69
DEMUCS [DEMUCS] 33.53 7.70 2.27 91.5 12.36 3.00 3.95 2.69 3.09 2.52 3.56 2.70 2.94 3.79 0.71
DeepFilterNet2 [DeepFilterNet2] 2.31 0.36 2.35 91.3 11.95 2.92 3.87 2.58 3.19 3.25 3.74 3.62 3.47 3.60 0.76
DeepFilterNet3 [DeepFilterNet3] 2.14 0.35 2.45 90.6 12.10 3.03 4.01 2.71 3.22 3.84 4.12 3.98 3.75 3.85 0.82
$\text{Baseline}^{*}$ 2.31 0.36 2.48 89.2 10.25 3.10 4.08 2.80 3.16 3.95 4.37 4.17 3.87 4.02 0.80
+OA Loss 2.31 0.36 2.56 90.2 11.21 3.13 4.06 2.82 3.20 3.80 4.28 4.07 3.79 4.00 0.84
   +Fine-Tuning 2.31 0.36 2.68 91.6 13.27 3.14 4.07 2.83 3.21 3.96 4.31 4.17 3.83 4.05 0.91
DPDFNet-2 2.49 1.35 2.79 92.6 13.72 3.16 4.08 2.85 3.25 4.06 4.36 4.23 3.92 4.10 0.95
DPDFNet-4 2.84 2.36 2.82 93.0 14.11 3.17 4.08 2.86 3.28 4.15 4.40 4.30 3.96 4.14 0.98
DPDFNet-8 3.54 4.37 2.85 93.4 14.47 3.19 4.09 2.89 3.28 4.21 4.43 4.34 4.00 4.18 1.00
@@ IV-A Implementation Details
We evaluate model performance using a combination of established intrusive and non-intrusive evaluation metrics. The intrusive metrics include PESQ [rix2001pesq], STOI [taal2011stoi], and SI-SNR [leroux2019sdr], which rely on comparisons between the enhanced signal and a clean, time-aligned reference. The non-intrusive metrics consist of DNSMOS P.835 (SIG, BAK, and OVRL [dnsmosp835]), P.808 MOS [dnsmosp808], and NISQA v2.0 [NISQA], which evaluates Overall Quality, Noisiness, Coloration, Discontinuity, and Loudness. Audio files were resampled to 16 kHz prior to processing. Training examples were created by segmenting clean speech into 3 sec chunks and mixing them with noise at SNRs randomly sampled from -5 dB to 40 dB. For STFT parameters, we followed the original DeepFilterNet2 setup, employing a window length of 20 ms, a hop size of 10 ms, and a Vorbis window function. We trained the models for 1.6M iterations with a batch size of 32, employing the AdamW [adamw] optimizer. The learning rate was decayed according to a cosine-annealing schedule, with a pick value of 1e-3.
Among the open-source, we select only those that meet two key criteria: (1) they are causal models suitable for streaming use, and (2) their weights are publicly available, allowing us to run the official versions on our evaluation set. Consequently, all models mentioned in Section I are included in our evaluations. For the models introduced in this work, we begin with a Baseline model that starts from the same architecture as DeepFilterNet2 [DeepFilterNet2], retrained within our framework for methodological consistency. We then present three variants of our proposed DPDFNet architecture: DPDFNet-2, DPDFNet-4, and DPDFNet-8. The suffix $k\in{2,4,8}$ indicates the number of DPRNN blocks in the dual-path stack, scaling the model’s capacity from small ($k=2$), to medium ($k=4$), and large ($k=8$).

### v2

@@ I Introduction
Early real-time speech enhancement often combined traditional signal processing with compact neural networks. Valin [26] proposed RNNoise, which predicts critical-band gains from features such as Bark-frequency cepstral coefficients and pitch correlations, then applies a frequency-domain pitch filter. This hybrid design achieves very low latency and embedded-class efficiency, and it outperforms minimum mean-square spectral estimators in perceptual quality. Research then advanced sequence modeling for streaming noise suppression. Westhausen and Meyer [27] introduced DTLN with two cascaded cores. One core operates in the frequency-domain and the other uses a learned time-domain representation. Both rely on LSTM layers with instant layer normalization, enabling stable frame-synchronous processing with fewer than one million parameters and benefiting from both magnitude and phase cues. Braun and Tashev proposed NSNet2 [3], a recurrent model that estimates spectral suppression gains from log power spectral features. Its level-invariant loss normalization and broad augmentation over SNR, spectral shape, and signal level yield a robust real-time baseline. Another line of work moved to the waveform domain. Defossez et al. [4] proposed DEMUCS, an encoder-decoder with U-Net-style skip connections and a recurrent bottleneck that operates directly on raw audio. Training uses a mix of waveform loss and multi-resolution spectral losses with strong augmentation. This approach improves naturalness and intelligibility under real-time constraints. Kong et al. proposed CleanUNet [8] as a refinement of this design. It strengthens the bottleneck with stacked masked self-attention and combines waveform and spectral objectives, including a high-band loss that improves the handling of silence and high-frequency detail. Although relatively large, CleanUNet maintains low latency and delivers consistent quality. Work in the time-frequency domain focused on stronger cross-band modeling while preserving local structure. Hao et al. [7] proposed FullSubNet, which encodes long-range context with a full-band LSTM. The frame-wise embedding is concatenated with local frequency neighborhoods and passed to a shared sub-band LSTM that predicts a complex ideal ratio mask [28] (cIRM) with tanh compression. This design links global spectral trends with local stationarity cues. Dual-path recurrent models were optimized to balance efficiency and streaming performance. Le et al. [10] proposed DPCRN, an encoder-decoder with causal two-dimensional convolutions and skip connections. A dual-path module models spectral structure within each frame and temporal dynamics across frames, under causal constraints. Rong et al. proposed GTCRN [20] as a direct follow-up. It uses grouped temporal convolutions and grouped dual-path recurrent units, merges high-frequency bands through an equivalent rectangular bandwidth (ERB) filter bank, and applies temporal recurrent attention. The result keeps the parameter count in the tens of thousands while remaining competitive with heavier systems. Recent work has also examined the impact of architectural simplicity on streaming efficiency. Ahn et al. [1] introduced FastEnhancer, which adopts a streamlined encoder-decoder design with the RNNFormer, combining recurrent temporal modeling with attention-based frequency modeling while preserving causality. By reducing memory operations and avoiding sequential assumptions over frequency, this approach highlights the importance of efficient axis-wise modeling for practical real-time speech enhancement. Most recently, Pei et al. proposed aTENNuate [15], a deep state space autoencoder for streaming enhancement on raw audio. The model stacks causal state space blocks with skip connections and light resampling. It targets online operation with steady latency and compact size, and it trains with objectives in both time and frequency domains. Overall, the field has made a substantial progress from classical-plus-RNN hybrids to highly capable pure end-to-end deep neural-network models, for higher intelligibility and naturalness under challenging real-time constraints.
@@
Although a wide array of methods developed in the recent years, including the aforementioned ones, we based our work over Schröter et al. [21]’s DeepFilterNet2 architecture: a compact two-stage architecture that offers a strong balance of efficiency and quality for streaming SE. This architecture provides a favorable trade-off between computational complexity and enhancement quality, enabling real-time, low-latency deployment. Moreover, the architecture permits training at lower sampling rates (e.g., 16 kHz) while deploying at higher rates with negligible performance loss. This leverages the wider availability of low-rate data and reduces the computational load of online augmentations. Motivated by both technical considerations and community interest in advanced speech enhancement, we believe that enhancing the current DeepFilterNet2 architecture will contribute much to the field.
@@
Accordingly, we integrate causal dual-path modules into the DeepFilterNet2 encoder, resulting in DPDFNet. This enhanced design delivers a substantial performance gain while preserving the overall architecture. We also add an over-attenuation loss, used alongside the original multi-resolution loss [21], to mitigate rare cases where the model over-suppresses the target speech, and proposed a fine-tuning stage to exposed the models to long-range dependencies. Finally, we curate a new evaluation set focused on low-SNR conditions with realistic noise types (car, pub, office, etc.) with multilingual coverage spanning 12 languages from the Speech-MASSIVE test set [11].
@@ II-C Dual-Path RNN Block
Dual-Path Recurrent Neural Networks, introduced by Luo et al. [13], address long-context modeling through two alternating recurrent stages that operate along complementary axes of the data: an intra stage that aggregates information within each time frame and an inter stage that propagates information across time.
@@
This dual-path strategy was adopted by DPCRN [10], which replaces the GRU bottleneck in a U-Net-style encoder-decoder with DPRNN blocks and reports strong performance. In addition, DPCRN appends a position-wise fully connected (FC) layer and Layer Normalization (LN) [2] after each DPRNN stage, both intra and inter, in order to refine features and stabilize training. Figure 1a depicts the scheme of the proposed DPRNN block.
@@ III-A Datasets and Augmentations
Many recent SE models, including our baseline DeepFilterNet2, have been trained on the widely adopted Deep Noise Suppression 4 (DNS4) Challenge [5] dataset. This full-band (48 kHz) dataset provides 760 hours of clean speech recordings in six languages from 3,230 unique speakers, along with 181 hours of noise comprising 62,000 clips across 150 noise classes. It also includes 248 real and 60,000 synthetic RIRs for generating reverberant signals. Although DNS4 remains a widely adopted dataset for training SE models, the clean speech portion is not entirely free of residual noise, which may constrain the achievable performance of trained models. In contrast, a variety of high-quality, open-source wideband (16 kHz) datasets exist for both clean speech and noise, offering valuable resources for more effective model training. Furthermore, since the DeepFilterNet2 architecture can be adapted to full-band operation via a lightweight postprocessing step with negligible impact on performance, utilizing these larger and higher-quality wideband datasets presents a particularly compelling direction for training. Therefore, we downsampled the DNS4 dataset to wideband and enriched it with 4,000 hours of English read speech from the Multilingual LibriSpeech (MLS) corpus [16], a large-scale dataset derived from LibriVox released by Facebook. This addition improves both quality and speaker diversity compared to DNS4 alone. To further ensure the cleanest possible speech signals, all audio was processed through the DPCRN [10] model. Furthermore, we incorporated noise clips from the MUSAN [23] and FSD50K [6] datasets to further diversify the noise conditions. To improve model generalization, a set of online data augmentations is applied during training. Random second-order filtering is used to expose the model to natural spectral colorations, improving robustness across devices and recording environments. Gain perturbations are introduced to prevent overfitting to specific loudness levels and to ensure invariance to overall volume. In addition, reverberation with early reflection targets is applied, enabling the model to suppress late reverberation while preserving the direct path and early reflections that are beneficial for speech intelligibility.
@@ III-B New Evaluation Set
To evaluate and compare different SE models, two test sets are commonly employed: (1) the VoiceBank+DEMAND test set [25] and (2) the DNS4 blind test set [5]. While these datasets have been widely adopted as benchmarks within the single-channel SE domain, they fail to fully capture the diversity of real-world acoustic conditions. For instance, the VoiceBank+DEMAND test set primarily consists of short clips, typically only a few seconds in duration, recorded at relatively high SNRs. Conversely, the DNS4 blind test set provides longer recordings; however, each clip includes only a single speaker and lacks extended periods of background noise without speech - an element often present in realistic environments. Furthermore, both datasets exhibit limited linguistic diversity, as they consist almost exclusively of English speech.
@@ III-C Loss Functions
We adopt the Multi-Resolution loss from  [21]. The enhanced signal $y$ is analyzed with several STFTs, specifically with window sizes $i\in\{5,10,20,40\}$ ms, to form the following loss:
@@
denote the complex spectrograms of the enhanced and clean signals, respectively, with phases $\phi_{Y_{i}}$ and $\phi_{S_{i}}$. The magnitude compression exponent set to $c=0.3$ follow the original DeepFilterNet2 training [21].
@@ IV Experimental Results
TABLE I: Intrusive, non-intrusive and PRISM scores on our new multilingual low-SNR evaluation set for open-source causal models and our DPDFNet-$\{k\}$ variants ($k$ = number of DPRNN blocks). The baseline uses the DeepFilterNet2 architecture within our training framework.
@@
Noisy 1.32 83.2 0.38 2.03 1.66 1.56 2.53 2.00 1.69 3.40 2.68 2.68 0.04
DTLN [27] 0.99 0.12 2.14 88.7 10.83 2.51 3.50 2.15 2.85 2.13 2.51 2.79 2.39 2.85 0.46
GTCRN [20] 0.023 0.039 2.21 87.2 9.11 2.56 3.76 2.25 3.00 2.53 3.27 3.17 2.80 3.01 0.49
RNNoise [26] 0.087 0.04 2.36 88.5 8.97 2.81 3.94 2.51 3.01 1.88 3.68 2.20 2.19 3.14 0.52
NSNet2 [3] 2.60 0.26 2.35 86.8 8.63 2.68 3.82 2.36 2.92 2.63 3.39 3.25 2.76 3.30 0.52
FastEnhancer-T [1] 0.022 0.055 2.39 87.6 10.75 2.60 3.91 2.32 2.98 2.74 3.84 3.09 2.95 3.44 0.58
FullSubNet [7] 5.60 30.00 2.36 89.6 10.63 2.91 3.39 2.43 3.06 2.76 2.78 3.83 3.41 3.29 0.63
DPCRN [10] 0.53 1.10 2.76 90.7 11.93 2.74 3.79 2.40 3.03 2.92 3.11 3.61 3.23 3.40 0.69
aTENNuate [15] 0.80 0.33 2.80 88.9 9.65 3.05 4.03 2.74 2.94 3.01 3.35 3.90 2.97 3.92 0.70
DEMUCS [4] 33.53 7.70 2.57 91.5 12.36 3.00 3.95 2.69 3.09 2.52 3.56 2.70 2.94 3.79 0.72
FastEnhancer-B [1] 0.092 0.26 2.63 89.6 11.88 2.82 3.99 2.53 3.09 3.40 4.13 3.67 3.44 3.76 0.73
DeepFilterNet2 [21] 2.31 0.36 2.59 91.2 11.95 2.93 3.87 2.58 3.19 3.24 3.74 3.62 3.47 3.60 0.75
CleanUNet [8] 46.07 15.44 2.82 93.0 12.50 3.08 4.00 2.77 3.15 2.83 3.65 2.96 3.03 3.71 0.79
FastEnhancer-M [1] 0.19 0.66 2.58 90.4 12.96 2.90 3.98 2.61 3.13 3.87 4.33 4.03 3.79 4.01 0.80
FastEnhancer-S [1] 0.49 2.90 2.75 90.4 12.51 2.89 4.03 2.61 3.15 3.71 4.32 3.94 3.67 3.97 0.81
DeepFilterNet3 [22] 2.14 0.35 2.76 90.5 12.10 3.03 4.01 2.72 3.22 3.84 4.12 3.99 3.75 3.85 0.82
FastEnhancer-L [1] 1.11 11.00 2.76 91.2 13.57 2.94 4.00 2.65 3.17 4.15 4.49 4.27 3.98 4.15 0.87
$\text{Baseline}^{*}$ 2.31 0.36 2.85 89.2 10.25 2.94 4.00 2.65 3.17 3.95 4.37 4.17 3.87 4.02 0.79
+OA Loss 2.31 0.36 2.92 90.2 11.21 3.19 4.09 2.89 3.28 3.80 4.28 4.07 3.79 4.00 0.85
   +Fine-Tuning 2.31 0.36 3.04 91.6 13.27 3.14 4.07 2.83 3.21 3.96 4.31 4.17 3.83 4.05 0.91
DPDFNet-2 2.49 1.35 3.14 92.6 13.72 3.16 4.08 2.85 3.26 4.06 4.36 4.23 3.92 4.10 0.95
DPDFNet-4 2.84 2.36 3.18 93.0 14.11 3.17 4.08 2.87 3.28 4.15 4.40 4.30 3.96 4.14 0.98
DPDFNet-8 3.54 4.37 3.20 93.4 14.47 3.19 4.09 2.89 3.28 4.21 4.43 4.34 4.00 4.18 1.00
@@ IV-A Implementation Details
We evaluate model performance using a combination of established intrusive and non-intrusive evaluation metrics. The intrusive metrics include PESQ [19], STOI [24], and SI-SNR [9], which rely on comparisons between the enhanced signal and a clean, time-aligned reference. The non-intrusive metrics consist of DNSMOS P.835 (SIG, BAK, and OVRL [18]), P.808 MOS [17], and NISQA v2.0 [14], which evaluates Overall Quality, Noisiness, Coloration, Discontinuity, and Loudness. Audio files were resampled to 16 kHz prior to processing. Training examples were created by segmenting clean speech into 3 sec chunks and mixing them with noise at SNRs randomly sampled from -5 dB to 40 dB. For STFT parameters, we followed the original DeepFilterNet2 setup, employing a window length of 20 ms, a hop size of 10 ms, and a Vorbis window function. We trained the models for 1.6M iterations with a batch size of 32, employing the AdamW [12] optimizer. The learning rate was decayed according to a cosine-annealing schedule, with a pick value of 1e-3.
Among the open-source, we select only those that meet two key criteria: (1) they are causal models suitable for streaming use, and (2) their weights are publicly available, allowing us to run the official versions on our evaluation set. Consequently, all models mentioned in Section I are included in our evaluations. For the models introduced in this work, we begin with a Baseline model that starts from the same architecture as DeepFilterNet2 [21], retrained within our framework for methodological consistency. We then present three variants of our proposed DPDFNet architecture: DPDFNet-2, DPDFNet-4, and DPDFNet-8. The suffix $k\in{2,4,8}$ indicates the number of DPRNN blocks in the dual-path stack, scaling the model’s capacity from small ($k=2$), to medium ($k=4$), and large ($k=8$).